# Progressively Selective Label Enhancement for Language Model Alignment

### v1

@@ Abstract
Large Language Models have demonstrated impressive capabilities in various language tasks but may produce content that misaligns with human expectations, raising ethical and legal concerns. Therefore, it is important to explore the limitations and implement restrictions on the models to ensure safety and compliance, with Reinforcement Learning from Human Feedback (RLHF) being the primary method. Due to challenges in stability and scalability with the RLHF stages, researchers are exploring alternative methods to achieve effects comparable to those of RLHF. However, these methods often depend on large high-quality datasets and inefficiently utilize generated data. To deal with this problem, we propose Psle, i.e., Progressively Selective Label Enhancement for Language Model Alignment, a framework that fully utilizes all generated data by guiding the model with principles to align outputs with human expectations. Using a dynamically updated threshold, our approach ensures efficient data utilization by incorporating all generated responses and weighting them based on their corresponding reward scores. Experimental results on multiple datasets demonstrate the effectiveness of Psle compared to existing language model alignment methods.
@@ 1 Introduction
However, these language models may not always generate text as expected by humans and can even produce content that violates human ethics or legal boundarie (Bai et al., 2022a; Askell et al., 2021). Therefore, it is important for researchers to explore the limitations of these models and implement restrictions on output generation to ensure safety and compliance, a process known as AI alignment.
@@
The most prominent method for achieving AI alignment is Reinforcement Learning from Human Feedback (RLHF) (Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022). RLHF employs Supervised Fine-Tuning (SFT) to guide models using human instructions (Wang et al., 2023b; Taori et al., 2023), followed by training a Reward Model on human-rated outputs (Ouyang et al., 2022), and optimizing the model with Reinforcement Learning (RL) algorithms like Proximal Policy Optimization (PPO) (Schulman et al., 2017; Askell et al., 2021; Bai et al., 2022b). However, due to challenges in stability and scalability with the RL stage, researchers are exploring alternative methods. For instances, LIMA (Zhou et al., 2023) has experimentally demonstrated that when the pre-trained model’s capabilities are sufficiently strong and the quality of the SFT data is high, it can achieve results comparable to those of RLHF. RAFT (Dong et al., 2023) expands the SFT dataset by generating additional samples and selecting those with high reward scores to enhance the SFT dataset. RRHF (Yuan et al., 2023) simplifies the RLHF process by integrating the subsequent RL steps into the SFT phase as a regularization term.
@@
However, these methods are highly dependent on large amounts of high-quality data, which is impractical in certain applications, such as the medical field (Yang et al., 2024; Li et al., 2023) or chip design (Liu et al., 2024, 2023). Additionally, even though some methods generate extra data to expand the training set to alleviate the problem. These methods often select only a small portion of high-scoring data from the reward model, discarding a significant amount of other potentially useful data, making the algorithms highly inefficient, since generation is time-consuming for language models. Therefore, we consider using a more efficient approach that fully utilizes all generated data, even if some of it may not score as highly. This data can still contribute to the model training process, ensuring that no valuable information is wasted and improving overall training efficiency.
@@
Motivated by the above consideration, we propose a novel framework named Psle, i.e., Progressively Selective Label Enhancement for Language Model Alignment. Specially, during the sample generation phase, we design a set of principles to guide the model to output according to human expectations. When the reward score difference between the principle-guided output and the original input’s response exceeds a dynamically updated threshold, indicating a significant improvement under the principle-guiding, the model is encouraged to align its output with the better response and move away from the poorer one. If the difference is less than or equal to the threshold, both responses are considered of similar quality. To fully utilize all generated responses, we involve both in the model’s training, assigning weights based on the normalized reward scores. Our contributions can be summarized as follows:
  • Practically, we are the first to identify the issue of low data utilization in existing methods and propose a new approach that effectively leverages all generated data, regardless of their scores, allowing for more efficient expansion of the dataset.
  • Theoretically, we prove that with the progressively updated threshold strategy, our approach can converge to the optimal language model, which ideally achieves the highest expected reward scores.
@@
Experimental results on multiple datasets validate the effectiveness of our method s over several existing language model alignment approaches.
@@ 2 Related Work
The alignment of language models refers to the process of ensuring that the models behave in ways that are consistent with human values, ethical principles, and intended purposes (Leike et al., 2018). The most prominent and effective methods currently used to achieve this alignment is Reinforcement Learning from Human Feedback (RLHF) (Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022). The framework of RLHF first employs Supervised Fine-Tuning (SFT) to guide the model in following human instructions with an imitative manner (Wang et al., 2023b; Taori et al., 2023). The next steps involve training a Reward Model on a dataset reflecting human preferences, created from human evaluators’ ratings of the SFT model’s outputs (Ouyang et al., 2022). Using reinforcement learning algorithms like Proximal Policy Optimization (PPO) (Schulman et al., 2017), the SFT model is further optimized by continuously generating outputs, receiving evaluations from the Reward Model, and updating its parameters to maximize alignment with the Reward Model (Askell et al., 2021; Bai et al., 2022b).
However, due to the challenges of stability and scalability involved in the interactions between multiple models in RLHF, researchers have started exploring other more direct and efficient methods for model alignment (Rafailov et al., 2023; Yuan et al., 2023; Zhou et al., 2023; Dong et al., 2023). DPO derives an equivalent optimization objective from RLHF, allowing the model to be directly optimized using preference data without the need to train a separate reward model (Rafailov et al., 2023). Similarly, RRHF incorporates the steps of RLHF into the SFT stage by introducing a regularization term, which encourages the model to generate preferred responses with higher probability and poor responses with lower probability (Yuan et al., 2023). LIMA has experimentally demonstrated that when the pre-trained model is sufficiently good, only a small amount of high-quality data is needed. By using just SFT, it is possible to obtain a well-aligned model without the need for the subsequent complex RLHF steps (Zhou et al., 2023). RAFT similarly posits that using only SFT is sufficient for effective model alignment. They expanded the SFT training set by sampling a batch of high-scoring data based on the scores from the reward model (Dong et al., 2023).
@@ 3 Preliminaries
We first introduce the formal notation for the language model alignment problem. Let $\mathcal{V}$ be a vocabulary of a language model. The goal of alignment is to ensure that the language model $\pi:\mathcal{X}\rightarrow\mathcal{Y}$ generates response $\bm{y}\in\mathcal{Y}$ that are consistent with human values and preferences given a query $\bm{x}\in\mathcal{X}$, where the query $\bm{x}=[x^{1},x^{2},\dots,x^{m}]$ and response $\bm{y}=[y^{1},y^{2},\dots,y^{n}]$ are sequences of tokens, the input space $\mathcal{X}=\mathcal{V}^{m}$ and the output space $\mathcal{Y}=\mathcal{V}^{n}$.
The alignment process typically begins with Supervised Fine-Tuning (SFT), which adjusts the language model using Maximum Likelihood Estimation on a human-labeled high-quality dataset $\mathcal{D}_{\text{sft}}=\{(\bm{x}_{i},\bm{y}_{i})\}_{i=1}^{N}$:
$\mathcal{L}_{\text{sft}}=-\sum_{i=1}^{N}\sum_{j=1}^{n_{i}}\log P(y_{i}^{j}|[y_{i}^{k}]_{k=0}^{j-1},\bm{x}^{(i)};\theta),$ (1)
where $N$ is the number of training examples, $n_{i}$ is the length of the $i$-th target sequence, and $\theta$ represents the parameters of the language model $\pi_{\theta}$. For the notational simplicity, $y_{i}^{0}=\emptyset$ denotes an empty placeholder.
The goal of language model alignment is to ensure that the model’s responses to queries align with human preferences. These preferences are typically captured by a reward model $R:(\mathcal{X},\mathcal{Y})\rightarrow\mathbb{R}$, where higher scores indicate responses that better align with human values and preferences. Conversely, lower scores indicate less alignment. An ideal model maximizes the expected reward:
$\pi^{*}=\arg\max_{\pi}\mathbb{E}_{\bm{x}\sim p(\bm{x}),\bm{y}\sim\pi(\bm{x})}[R(\bm{x},\bm{y})],$ (2)
where $\pi^{*}$ represents the optimal policy that maximizes the expected reward according to the reward model $R$.
@@ 4 The Proposed Method
Figure 1: Comparison of language model responses with and without principle guidance. (a) Without principles, the model provides an unethical response to a query about embezzlement. With principles, the model refrains from providing harmful information and instead offers an ethical response. (b) For a query about job interview attire, both responses with principle guidance are consistent and align with being informative and helpful, showing improved quality and alignment with human values.
Algorithm 1 The Psle Algorithm
@@
3:  Initialize the training dataset with a empty set $\mathcal{D}_{\text{train}}=\emptyset$
@@
7:        Sample a response $\bm{y}\sim\pi_{\theta}(\bm{x})$
8:        Sample a principle-guided response $\bm{y}^{\text{prompt}}\sim\pi_{\theta}([\bm{p},\bm{x}])$
@@
14:        Update threshold $\tau=\tau_{0}\cdot\alpha^{t}$
15:     end for
@@
In this section, we present our novel framework named Psle, i.e., Progressively Selective Label Enhancement for Language Model Alignment. As illustrated in Figure 1, during the sample generation phase, we use principles to direct the model’s outputs. When the reward score difference between the principle-guided output and the original response exceeds a dynamically updated threshold, the model is encouraged to align with the better response and move away from the poorer one. If the difference is below the threshold, both responses are used in training since both are considered of similar quality.
4.1 Psle
Language model alignment requires a large amount of high-quality data, but this is impractical in many scenarios. Therefore, we consider generating additional data during training to expand the dataset. Motivated by the self-align approach (Wang et al., 2023b; Sun et al., 2023), we design a set of principles to guide the model in generating responses that align closely with human preferences:
@@ 4.1 Psle
Let $\pi^{\text{sft}}_{\theta}$ be the SFT-aligned model optimized by Eq. (1) and we use it as the initial model. During training, for each query $\bm{x}\in\mathcal{D}_{\text{query}}=\{\bm{x}_{i}\}_{i=1}^{N_{q}}$, where $N_{q}$ is the number of queries, the model samples a response $\bm{y}\sim\pi_{\theta}(\bm{x})$. In addition, the principle-guided model then samples a response: $\bm{y}^{\text{prompt}}\sim\pi_{\theta}([\bm{p},\bm{x}])$ based on the set of principles designed to ensure ethical, informative, and helpful output. The reward model $R$ assigns the scores $s=R(\bm{x},\bm{y})$ and $s^{\text{prompt}}=R(\bm{x},\bm{y}^{\text{prompt}})$.
@@
$\mathcal{L}_{\text{rank}}=\sum_{{s^{\text{prompt}}-s>\tau}}\pi_{\theta}(\bm{y}|\bm{x})-\pi_{\theta}(\bm{y}^{\text{prompt}}|\bm{x}).$ (3)
When the difference between the reward scores, $s^{\text{prompt}}-s$, is less than or equal to the threshold $\tau$, we consider that the response generated by the principle-guided model and the original response are of similar quality. Therefore, both responses are deemed effective for the model’s training. We include both responses in the dataset for subsequent training, with their weights determined by the magnitude of their scores. The formula for this process is as follow:
$\mathcal{L}_{\text{weighted-sft}}=-\sum_{s^{\text{prompt}}-s\leq\tau}\left(w\cdot\pi_{\theta}(\bm{y}|\bm{x})+w^{\text{prompt}}\cdot\pi_{\theta}(\bm{y}^{\text{prompt}}|\bm{x})\right),$ (4)
where the weights $w$ and $w^{\text{prompt}}$ are calculated as:
$w=\frac{e^{s}}{e^{s}+e^{s^{\text{prompt}}}},\quad w^{\text{prompt}}=\frac{e^{s^{\text{prompt}}}}{e^{s}+e^{s^{\text{prompt}}}}.$ (5)
This approach ensures that both the original and the principal-guided responses contribute to the training process, with their influence proportional to their respective reward scores. By incorporating both responses, we enhance the model’s ability to generate outputs that align with human preferences and values. Then the final objective function is:
@@
The whole process of Psle is shown in Algorithm 1.

### v2

# Progressively Label Enhancement for Large Language Model Alignment

@@ Abstract
Large Language Models (LLM) alignment aims to prevent models from producing content that misaligns with human expectations, which can lead to ethical and legal concerns. In the last few years, Reinforcement Learning from Human Feedback (RLHF) has been the most prominent method for achieving alignment. Due to challenges in stability and scalability with RLHF stages, which arise from the complex interactions between multiple models, researchers are exploring alternative methods to achieve effects comparable to those of RLHF. However, these methods often rely on large high-quality datasets. Despite some methods considering the generation of additional data to expand datasets, they often treat model training and data generation as separate and static processes, overlooking the fact that these processes are highly interdependent, leading to inefficient utilization of the generated data. To deal with this problem, we propose Ple, i.e., Progressively Label Enhancement for LLM Alignment, a framework that dynamically adjusts the model’s training process based on the evolving quality of the generated data. Specifically, we prompt the model to generate responses for both the original query and the query guided by a set of carefully designed principles, and then utilize a dynamic threshold to determine the appropriate training approach for both responses based on their corresponding reward scores. Experimental results demonstrate the effectiveness of Ple compared to existing LLM alignment methods.
@@ 1 Introduction
However, these language models may not always generate responses as expected by humans and can even produce content that violates human ethics or legal boundaries (Bai et al., 2022a; Askell et al., 2021). Therefore, it is crucial for researchers to explore the limitations of these models and implement restrictions on output generation to ensure safety and compliance, a process known as AI alignment.
@@
The most prominent method for achieving AI alignment is Reinforcement Learning from Human Feedback (RLHF) (Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022). RLHF employs Supervised Fine-Tuning (SFT) to guide models using human instructions (Wang et al., 2023b; Taori et al., 2023), followed by training a reward model on human-rated outputs (Ouyang et al., 2022), and optimizing the model with Reinforcement Learning (RL) algorithms like Proximal Policy Optimization (PPO) (Schulman et al., 2017; Askell et al., 2021; Bai et al., 2022b). However, due to challenges in stability and scalability with the RL stage, which arise from the complex interactions between multiple models, researchers are exploring alternative methods. For instance, LIMA (Zhou et al., 2023) has experimentally demonstrated that when the pre-trained model’s capabilities are sufficiently strong and the quality of the SFT data is high, it can achieve results comparable to those of RLHF. RAFT (Dong et al., 2023) expands the SFT dataset by generating additional samples and selecting those with high reward scores to enhance the SFT dataset. RRHF (Yuan et al., 2023) simplifies the RLHF process by integrating the subsequent RL steps into the SFT phase as a regularization term.
@@
However, these methods are highly dependent on large amounts of high-quality data, which is impractical in certain applications, such as the medical field (Yang et al., 2024; Li et al., 2023) or chip design (Liu et al., 2023). Additionally, even though some methods generate extra data to expand the training set to alleviate the problem. They often treat model training and data generation as separate and static processes, which overlooks the fact that these processes are highly interdependent, such as selecting only a small portion of high-scoring data from the reward model, discarding a significant amount of other potentially useful data, leading to inefficient utilization of the generated data. Therefore, we consider designing an efficient framework that couples the data generation and model training processes, allowing them to work synergistically, thus ensuring that all generated data, including potentially useful lower-scoring data, is effectively utilized, thereby improving training efficiency.
@@
Motivated by the above consideration, we propose a novel framework named Ple, i.e., Progressively Label Enhancement for Language Model Alignment. Specifically, during the sample generation phase, we design a set of principles to guide the model to output according to human expectations. When the reward score difference between the principle-guided output and the response to the original query exceeds a dynamically updated threshold, indicating a significant improvement under the principle guiding, the model is encouraged to align its output with the better response and move away from the poorer one. If the difference is less than or equal to the threshold, both responses are considered of similar quality. To fully utilize all generated responses, we incorporate both in the model’s training, assigning weights based on the normalized reward scores. Our contributions can be summarized as follows:
  • Practically, we are the first to identify that previous alignment methods overlook the coupling between data generation and model training, leading to inefficient utilization of generated data. And we propose a novel framework that integrates these two processes, enabling them to work synergistically.
  • Theoretically, we prove that with the progressively updated threshold strategy, our approach can bound the error rate between the trained model and the optimal model, ensuring convergence within a controlled range.
@@
Extensive experimental results validate the effectiveness of our methods over several existing language model alignment approaches.
@@ 2 Related Work
The alignment of language models refers to the process of ensuring that the models behave in ways that are consistent with human values, ethical principles, and intended purposes (Leike et al., 2018). The most prominent and effective method currently used to achieve this alignment is Reinforcement Learning from Human Feedback (RLHF) (Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022). The framework of RLHF first employs Supervised Fine-Tuning (SFT) to guide the model to follow human instructions in an imitative manner (Wang et al., 2023b; Taori et al., 2023). The next steps involve training a Reward Model on a dataset reflecting human preferences, created from human evaluators’ ratings of the SFT model’s outputs (Ouyang et al., 2022). Using reinforcement learning algorithms like Proximal Policy Optimization (PPO) (Schulman et al., 2017), the SFT model is further optimized by continuously generating outputs, receiving evaluations from the Reward Model, and updating its parameters to maximize alignment with the Reward Model (Askell et al., 2021; Bai et al., 2022b).
However, due to the challenges of stability and scalability involved in the interactions between multiple models in RLHF, researchers have started exploring other more direct and efficient methods for model alignment (Rafailov et al., 2023; Yuan et al., 2023; Zhou et al., 2023; Dong et al., 2023). DPO derives an equivalent optimization objective from RLHF, allowing the model to be directly optimized using preference data without the need to train a separate reward model (Rafailov et al., 2023). Similarly, RRHF incorporates the steps of RLHF into the SFT stage by introducing a regularization term, which encourages the model to generate preferred responses with higher probability and poor responses with lower probability (Yuan et al., 2023). LIMA has experimentally demonstrated that when the pre-trained model is sufficiently good, only a small amount of high-quality data is needed. By using only SFT, it is possible to obtain a well-aligned model without the need for the subsequent complex RLHF steps (Zhou et al., 2023). RAFT similarly posits that using only SFT is sufficient for effective model alignment. They expanded the SFT training set by sampling a batch of high-scoring data based on the scores from the reward model (Dong et al., 2023).
@@ 3 Preliminaries
We first introduce the formal notation for the language model alignment problem. Let $\mathcal{V}$ be a vocabulary of a language model. The goal of alignment is to ensure that the language model $\pi:\mathcal{X}\rightarrow\mathcal{Y}$ generates response $\bm{y}\in\mathcal{Y}$ that is consistent with human values and preferences given a query $\bm{x}\in\mathcal{X}$, where the query $\bm{x}=[x^{1},x^{2},\dots,x^{m}]$ and response $\bm{y}=[y^{1},y^{2},\dots,y^{n}]$ are sequences of tokens, the input space $\mathcal{X}=\mathcal{V}^{m}$ and the output space $\mathcal{Y}=\mathcal{V}^{n}$.
The alignment process typically begins with Supervised Fine-Tuning (SFT) stage, which adjusts the language model using Maximum Likelihood Estimation on a human-labeled high-quality dataset $\mathcal{D}_{\text{sft}}=\{(\bm{x}_{i},\bm{y}_{i})\}_{i=1}^{N}$:
$\mathcal{L}_{\text{sft}}=-\sum_{i=1}^{N}\sum_{j=1}^{n_{i}}\log P(y_{i}^{j}|[y_{i}^{k}]^{k<j},\bm{x}_{i};\theta),$ (1)
where $N$ is the number of training examples, $n_{i}$ is the length of the $i$-th target sequence, and $\theta$ represents the parameters of the language model $\pi_{\theta}$.
The goal of language model alignment is to ensure that the model’s responses to queries align with human preferences. These preferences are typically captured by a reward model $R:(\mathcal{X},\mathcal{Y})\rightarrow\mathbb{R}$, where higher scores indicate that responses better align with human values and preferences. Conversely, lower scores indicate less alignment. An ideal model maximizes the expected reward:
$\pi^{\star}=\arg\max_{\pi}\mathbb{E}_{\bm{x}\sim p(\bm{x}),\bm{y}\sim\pi(\cdot|\bm{x})}[R(\bm{x},\bm{y})],$ (2)
where $\pi^{\star}$ represents the optimal policy that maximizes the expected reward according to the reward model $R$.
@@ 4 The Proposed Method
Figure 1: Comparison of language model responses with and without principle guidance. (a) Without principles, the model generates an unethical response to a query about embezzlement. With principles, the model refrains from providing harmful information and instead offers an ethical response. (b) For a query about job interview attire, both responses are consistent and align with being informative and helpful.
Algorithm 1 The Ple Algorithm
@@
3:  Initialize the training dataset with a empty set $\mathcal{D}_{\text{train}}=\emptyset$.
@@
7:        Sample a response $\bm{y}\sim\pi_{\theta}(\cdot|\bm{x})$
8:        Sample a principle-guided response $\bm{y}^{\text{prompt}}\sim\pi_{\theta}(\cdot|[\bm{p},\bm{x}])$
@@
14:     end for
15:     Update threshold $\tau=\tau_{0}\cdot\alpha$
@@
In this section, we present our novel framework named Ple, i.e., Selective Label Enhancement for Language Model Alignment. As illustrated in Figure 1, during the sample generation phase, we use carefully crafted principles to guide the model’s outputs. When the reward score difference between the principle-guided output and the original response exceeds a dynamically updated threshold, the model is encouraged to align with the better response and move away from the poorer one. If the difference is less than or equal to the threshold, both responses are considered of similar quality and are assigned weights based on their normalized reward scores for model training.
4.1 Ple
Language model alignment requires a large amount of high-quality data, which is often impractical in many scenarios. Therefore, we consider generating additional data during training to expand the dataset. Motivated by the self-align approach (Wang et al., 2023b; Sun et al., 2023), we design a set of principles to guide the model in generating responses that align closely with human preferences:
@@ 4.1 Ple
Let $\pi^{\text{sft}}_{\theta}$ be the SFT-aligned model optimized by Eq. (1) and we use it as the initial model. During training, for each query $\bm{x}\in\mathcal{D}_{\text{query}}=\{\bm{x}_{i}\}_{i=1}^{N_{q}}$, where $N_{q}$ is the number of queries, the model samples a response $\bm{y}\sim\pi_{\theta}(\cdot|\bm{x})$. In addition, the principle-guided model then samples a response: $\bm{y}^{\text{prompt}}\sim\pi_{\theta}(\cdot|[\bm{p},\bm{x}])$ based on the set of principles designed to expect ethical, informative, and helpful output. The reward model $R$ assigns the scores $s=R(\bm{x},\bm{y})$ and $s^{\text{prompt}}=R(\bm{x},\bm{y}^{\text{prompt}})$.
@@
$\mathcal{L}_{\text{rank}}=-\sum_{{s^{\text{prompt}}-s>\tau}}\pi_{\theta}(\bm{y}^{\text{prompt}}|\bm{x})-\pi_{\theta}(\bm{y}|\bm{x}).$ (3)
When the difference between the reward scores is less than or equal to the threshold $\tau$, we consider that the response generated by the principle-guided model and the original response are of similar quality. Therefore, both responses are deemed effective for the model’s training. We include both responses in the dataset for subsequent training, with their weights determined by the magnitude of their scores, a process known as label enhancement (Xu et al., 2021; 2023b). The formula for this process is as follows:
$\mathcal{L}_{\text{weighted-sft}}=-\sum_{s^{\text{prompt}}-s\leq\tau}w\cdot\pi_{\theta}(\bm{y}|\bm{x})+w^{\text{prompt}}\cdot\pi_{\theta}(\bm{y}^{\text{prompt}}|\bm{x}),$ (4)
where the weights $w$ and $w^{\text{prompt}}$ are calculated as follows to normalize them to the range $[-1,1]$:
$w=\frac{2e^{s}}{e^{s}+e^{s^{\text{prompt}}}}-1,\quad w^{\text{prompt}}=\frac{2e^{s^{\text{prompt}}}}{e^{s}+e^{s^{\text{prompt}}}}-1.$ (5)
This approach ensures that both the original and the principle-guided responses contribute to the training process, with their influence proportional to their respective reward scores. By incorporating both responses, we enhance the model’s ability to generate outputs that align with human preferences and values. Then the final objective function is:
@@
The whole process of Ple is shown in Algorithm 1.
5 Theoretical Analysis
In this section, we will provide a theoretical analysis to demonstrate that Ple, which uses a dynamically updated threshold for data selection and model training, will ultimately converge to the optimal model $\pi^{\star}$ defined in Eq. (2).
Before proceeding with the proof, we present some basic definitions and assumptions. For two queries $\bm{x}$ and $\bm{z}$ that satisfy $R(\bm{z},\bm{y}_{\bm{z}}^{\text{prompt}})-R(\bm{z},\bm{y}_{\bm{z}})>R(\bm{x},\bm{y}_{\bm{x}}^{\text{prompt}})-R(\bm{x},\bm{y}_{\bm{x}})$, i.e., the margin between the reward score of principle-guided response $R(\bm{z},\bm{y}_{\bm{z}}^{\text{prompt}})$ and that of original response $R(\bm{z},\bm{y}_{\bm{z}})$ is larger than that in point $\bm{x}$, the indicator function $\Big{[}\bm{1}_{\{\pi(\bm{y}_{\bm{z}}^{\text{prompt}}|\bm{z})<\pi(\bm{y}_{\bm{z}}|\bm{z})\}}\Big{|}R(\bm{z},\bm{y}_{\bm{z}}^{\text{prompt}})-R(\bm{z},\bm{y}_{\bm{z}})>R(\bm{x},\bm{y}_{\bm{x}}^{\text{prompt}})-R(\bm{x},\bm{y}_{\bm{x}})\Big{]}$ equals 1 if the model’s output probabilities for the responses $\pi(\bm{y}_{\bm{z}}|\bm{z})$ and $\pi(\bm{y}_{\bm{z}}^{\text{prompt}}|\bm{z})$ are inconsistent with the corresponding ranking of their reward scores. Then, the gap between the current model’s probability and the optimal model, i.e., the approximation error of the model, could be controlled by the inconsistency between the model’s output probabilities and the corresponding ranking of their reward scores for all the queries $\bm{z}$.
Therefore, we assume that there exist constants $\alpha,\epsilon<1$, such that for any $\bm{x}\in\mathcal{X}$ and $\bm{y}\in\mathcal{Y}$,
$|\pi(\bm{y}|\bm{x})-\pi^{\star}(\bm{y}|\bm{x})|\leq\alpha\mathbb{E}_{(\bm{z},\bm{y},\bm{y}^{\text{prompt}})\sim\mathcal{D}_{\text{train}}}\Big{[}\bm{1}_{\{\pi(\bm{y}_{\bm{z}}^{\text{prompt}}|\bm{z})<\pi(\bm{y}_{\bm{z}}|\bm{z})\}}\Big{|}$ (8)
$R(\bm{z},\bm{y}_{\bm{z}}^{\text{prompt}})-R(\bm{z},\bm{y}_{\bm{z}})>R(\bm{x},\bm{y}_{\bm{x}}^{\text{prompt}})-R(\bm{x},\bm{y}_{\bm{x}})\Big{]}+\frac{\epsilon}{6}.$
In addition, for the probability density function $d(u)$ of the cumulative distribution function of the margin of the reward scores $D(u)=P_{\bm{x}\sim p(\bm{x})}(u(\bm{x})\leq u)$, where $u(\bm{x})=R(\bm{x},\bm{y}_{\bm{x}}^{\text{prompt}})-R(\bm{x},\bm{y}_{\bm{x}})$ denotes the margin of the reward scores for the query $\bm{x}$. We assume that there exists constants $c_{\star},c^{\star}$, such that $c_{\star}<d(u)<c^{\star}$. Then, we define the worst-case density imbalance ratio as $l=\frac{c^{\star}}{c_{\star}}$.
Motivated by the pure level set in traditional machine learning (Zhang et al., 2021; Xu et al., 2023a), the region where the language model is reliable, i.e., the region where the model’s output probabilities are consistent with the ranking of the reward scores of the principle-guided response and the original response, can be defined as:
Definition 5.1.
Pure ($e$, $\pi$, $R$)-level set: A queries set $L(e,R):=\{\bm{x}|R(\bm{x},\bm{y}_{\bm{x}}^{\text{prompt}})-R(\bm{x},\bm{y}_{\bm{x}})\geq e\}$ is pure for the model $\pi$, if for any $\bm{x}\in L(e,R)$, $\pi(\bm{y}_{\bm{x}}^{\text{prompt}}|\bm{x})>\pi(\bm{y}_{\bm{x}}|\bm{x})$.
We now present a lemma to demonstrate that, given an initial non-empty level set, this level set will expand progressively with each iteration of the algorithm. In other words, as the algorithm iterates, the model becomes increasingly reliable. Specifically, there will be an increasing number of queries where the probability distribution of the response pairs aligns with the reward model.
Lemma 5.2.
For a given language model $\pi$, there exists a pure $L(e,\pi,R)$-level set. For query $\bm{x}\in\mathcal{D}_{\text{query}}$, if $\pi(\bm{y}^{\text{prompt}}|\bm{x})-\pi(\bm{y}|\bm{x})>e$, we add the instance-responses pair into the preference dataset $\mathcal{D}_{\text{train}}$ for calculating ranking loss. And assume the updated model $\pi_{\text{new}}=\arg\min\mathbb{E}_{(\bm{x},\bm{y}^{\text{prompt},\bm{y}})\sim\mathcal{D}_{\text{train}}}\mathcal{L}_{\text{rank}}(\bm{x},\bm{y}^{\text{prompt}},\bm{y})$. Let $e_{\text{new}}=\min\{e|e>0,L(e,R)\text{ is pure for }\pi_{\text{new}}\}$ and assume that $e_{\text{new}}>\epsilon$ Then,
$R(\bm{y}^{\text{prompt}}|\bm{x})-e_{\text{new}}\geq(1+\frac{\epsilon}{6\alpha l})(R(\bm{y}^{\text{prompt}}|\bm{x})-e).$
The detail of the proof is provided in Appendix A.1. Lemma 5.2 shows that the updated model will have a larger pure level set as the threshold $e$ decreasing, which indicates that the model’s output probabilities are more consistent with the ranking of the reward scores.
Finally, we present the main theorem to demonstrate that the Ple algorithm will bound the difference between the learned model and the optimal model $\pi^{\star}$, provided there exists a pure level set for the initialized model.
Theorem 5.3.
Suppose there exists a pure $L(e_{0},\pi_{0},R)$-level set for the initialized model $\pi_{0}$, if one runs purification in the Ple algorithm with enough iterations and the initialization: (1) $e_{0}\geq\frac{\alpha+\frac{\epsilon}{6}}{1+\alpha}$, (2) $e_{\text{end}}>\epsilon$ (3) The iteration steps $I\geq\frac{6l}{\epsilon}\log(\frac{1-\epsilon}{\frac{1}{|\mathcal{Y}|}-e_{0}})$ , then we have:
$\mathbb{P}_{\bm{x}\sim p(\bm{x}),\bm{y}\sim p(\bm{y})}\big{(}|\pi(\bm{y}|\bm{x})-\pi^{\star}(\bm{y}|\bm{x})|>\frac{\epsilon}{2}\big{)}\leq 1-c_{\star}\epsilon.$ (9)
The proof of Theorem 5.3 is provided in Appendix A.2. This result provides theoretical support for the effectiveness of our method in aligning language models with generated preference data.